# Collaborative Perceiver: Elevating Vision-based 3D Object Detection via Local Density-Aware Spatial Occupancy

### v1

@@ Abstract
Vision-based bird’s-eye-view (BEV) 3D object detection has advanced significantly in autonomous driving by offering cost-effectiveness and rich contextual information. However, existing methods often construct BEV representations by collapsing extracted object features, neglecting intrinsic environmental contexts, such as roads and pavements. This hinders detectors from comprehensively perceiving the characteristics of the physical world. To alleviate this, we introduce a multi-task learning framework, Collaborative Perceiver (CoP), that leverages spatial occupancy as auxiliary information to mine consistent structural and conceptual similarities shared between 3D object detection and occupancy prediction tasks, bridging gaps in spatial representations and feature refinement. To this end, we first propose a pipeline to generate dense occupancy ground truths incorporating local density information (LDO) for reconstructing detailed environmental information. Next, we employ a voxel-height-guided sampling (VHS) strategy to distill fine-grained local features according to distinct object properties. Furthermore, we develop a global-local collaborative feature fusion (CFF) module that seamlessly integrates complementary knowledge between both tasks, thus composing more robust BEV representations. Extensive experiments on the nuScenes benchmark demonstrate that CoP outperforms existing vision-based frameworks, achieving 49.5% mAP and 59.2% NDS on the test set. Code and supplementary materials are available at this link.
@@ III-C Voxel-Height-guided Sampling
Method Input Size NDS $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$ mAP $\uparrow$
FCOS3D [wang2021fcos3d] 1600 $\times$ 900 42.8 0.690 0.249 0.452 1.434 0.124 35.8
DETR3D [wang2022detr3d] 1600 $\times$ 900 47.9 0.641 0.255 0.394 0.845 0.133 41.2
Ego3RT [lu2022learning] 1500 $\times$ 900 47.3 0.549 0.264 0.433 1.014 0.145 42.5
BEVFormer [li2022bevformer] 1600 $\times$ 900 53.5 0.631 0.257 0.405 0.435 0.143 44.5
CMT-C [yan2023cross] 1600 $\times$ 640 48.1 0.616 0.248 0.415 0.904 0.147 42.9
BEVDet [huang2021bevdet] 1600 $\times$ 640 48.2 0.529 0.236 0.395 0.979 0.152 42.2
PETRv2 [liu2023petrv2] 1600 $\times$ 640 55.3 0.601 0.249 0.391 0.382 0.123 45.6
X3KD [klingner2023x3kd] 1408 $\times$ 512 56.1 0.506 0.253 0.414 0.366 0.131 45.6
SOGDet [zhou2024sogdet] 1408 $\times$ 512 58.1 0.471 0.246 0.389 0.330 0.128 47.4
CoP (Ours) 1408 $\times$ 512 59.2 0.482 0.244 0.385 0.314 0.131 49.5
@@ III-D Global-Local Collaborative Feature Fusion
Based on the extracted local and global features, we further employ a global-local collaborative feature fusion (CFF) strategy to achieve effective feature refinement, enhancement, and interaction. It leverages the synergistic cues from both extracted features and provides a seamless integration of complementary knowledge across tasks. To this end, we introduce parallel pathways to fuse the local $f_{l}$ and global $f_{g}$ features. As depicted in Figure 4(b), our core concept involves selecting and combining the most relevant spatial and structural features from two heterogeneous, compressed feature sets. Initially, we incorporate average- and max-pooling operations, followed by two linear layers, to refine contextual information as described in Equation 6:

### v3

@@ Abstract
Vision-based bird’s-eye-view (BEV) 3D object detection has advanced significantly in autonomous driving by offering cost-effectiveness and rich contextual information. However, existing methods often construct BEV representations by collapsing extracted object features, neglecting intrinsic environmental contexts, such as roads and pavements. This hinders detectors from comprehensively perceiving the characteristics of the physical world. To alleviate this, we introduce a multi-task learning framework, Collaborative Perceiver (CoP), that leverages spatial occupancy as auxiliary information to mine consistent structural and conceptual similarities shared between 3D object detection and occupancy prediction tasks, bridging gaps in spatial representations and feature refinement. To this end, we first propose a pipeline to generate dense occupancy ground truths incorporating local density information (LDO) for reconstructing detailed environmental information. Next, we employ a voxel-height-guided sampling (VHS) strategy to distill fine-grained local features according to distinct object properties. Furthermore, we develop a global-local collaborative feature fusion (CFF) module that seamlessly integrates complementary knowledge between both tasks, thus composing more robust BEV representations. Extensive experiments on the nuScenes benchmark demonstrate that CoP outperforms existing vision-based frameworks, achieving 49.5% mAP and 59.2% NDS on the test set. Code is available at this link.
@@ III-C Voxel-Height-guided Sampling
Method Modality NDS $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$ mAP $\uparrow$
InfoFocus [wang2020infofocus] Lidar 39.5 0.363 0.265 1.132 1.000 0.395 39.5
PointPillars [lang2019pointpillars] Lidar 45.3 0.517 0.290 0.500 0.316 0.368 30.5
CenterFusion [nabati2021centerfusion] Lidar & Radar 44.9 0.326 0.631 0.261 0.516 0.614 32.6
MVFusion [wu2023mvfusion] Lidar & Radar 51.7 0.569 0.246 0.379 0.781 0.128 45.3
FCOS3D [wang2021fcos3d] Camera 42.8 0.690 0.249 0.452 1.434 0.124 35.8
DETR3D [wang2022detr3d] Camera 47.9 0.641 0.255 0.394 0.845 0.133 41.2
Ego3RT [lu2022learning] Camera 47.3 0.549 0.264 0.433 1.014 0.145 42.5
BEVFormer [li2022bevformer] Camera 53.5 0.631 0.257 0.405 0.435 0.143 44.5
CMT-C [yan2023cross] Camera 48.1 0.616 0.248 0.415 0.904 0.147 42.9
BEVDet [huang2021bevdet] Camera 48.2 0.529 0.236 0.395 0.979 0.152 42.2
PETRv2 [liu2023petrv2] Camera 55.3 0.601 0.249 0.391 0.382 0.123 45.6
X3KD [klingner2023x3kd] Camera 56.1 0.506 0.253 0.414 0.366 0.131 45.6
SOGDet [zhou2024sogdet] Camera 58.1 0.471 0.246 0.389 0.330 0.128 47.4
CoP (Ours) Camera 59.2 0.482 0.244 0.385 0.314 0.131 49.5
@@ III-D Global-Local Collaborative Feature Fusion
Based on the extracted local and global features and inspired by [shi2024cobev, hou2024fastocc, yu2023flashocc], we further employ a global-local collaborative feature fusion (CFF) strategy to achieve effective feature refinement, enhancement, and interaction. It leverages the synergistic cues from both extracted features and provides a seamless integration of complementary knowledge across tasks. To this end, we introduce parallel pathways to fuse the local $f_{l}$ and global $f_{g}$ features. As depicted in Figure 4(b), our core concept involves selecting and combining the most relevant spatial and structural features from two heterogeneous, compressed feature sets. Initially, we incorporate average- and max-pooling operations, followed by two linear layers, to refine contextual information as described in Equation 6: